OpenStudy (anonymous):

Rationalize the denominator. index of 3√5/√4)

OpenStudy (anonymous):

i can write this out if you want me to, i know its hard to understand

OpenStudy (anonymous):

\[\sqrt[3]{\frac{5}{4}}\]?

OpenStudy (anonymous):

yea if thats an index of 3, its kinda small

OpenStudy (anonymous):

then multiply top and bottom by \[\sqrt[3]{2}\] to get \[\frac{\sqrt[3]{10}}{2}\]

OpenStudy (anonymous):

i just have no clue what to do. how does the index of 3 affect it? am i supposed to multiply the numer and denom by radical 4?

OpenStudy (anonymous):

no just \[\sqrt[3]{2}\] because \[\sqrt[3]{8}=2\]

OpenStudy (anonymous):

where did you get the the radical 2? it went from radical 4 to radical 2?

OpenStudy (anonymous):

it is not 'radical 4' it is 'cubed root of 4'

OpenStudy (anonymous):

ok so first step was simplify the 4 into a 2, then cube that number?

OpenStudy (anonymous):

so it goes like this \[\sqrt[3]{4}\times \sqrt[3]{2}=\sqrt[3]{8}=2\] because \[2^3=8\] that gets rid of the cubed root in the denominator

OpenStudy (anonymous):

so how did you know to mult the 4 by 2? is it because its cubed?

OpenStudy (anonymous):

no i did not "simplify the 4" it is the cubed root of 4. i wanted to multiply it by something to give a cube

OpenStudy (anonymous):

so for example is the denom was 16, what would you do?

OpenStudy (anonymous):

so for example if i had \[\sqrt[3]{7}\] i would multiply by \[\sqrt[3]{49}\] to get \[\sqrt[3]{7^3}=7\]

OpenStudy (anonymous):

\[\sqrt[3]{5/\sqrt{16}}\]

OpenStudy (anonymous):

and if i had \[\sqrt[3]{16}=\sqrt[3]{2^4}\] i would use \[\sqrt[3]{2^2}\] to give \[\sqrt[3]{2^6}=2^3\]